A local Bank has a branch at Cape College. The branch is busiest at the beginning of the college year when new students open accounts. This year, students arrived at the office at a rate of 40 per day (8-hour day). On average, it takes the local Bank staff person about ten minutes to process each account application. The bank is considering having one or two tellers. Each teller is paid R120 per hour and the cost of waiting in line is assumed to be R80 per hour.

 

1. What is the total daily service cost for the single teller model? Show all steps!

2. What is the total daily service cost for the two-teller model? Show all steps!

3. Which model is preferred? Why?


1
Expert's answer
2021-03-16T09:08:39-0400

 

i)       Single teller

Arrival rate = 40 per day Service rate = 8*60/10 = 48 per day

Salary

R120 per hour

For 8hours

R120*8 = R960


Waiting

1 student =10min

40students = 40min

R80 waiting cost per hour

R80*40min/60min= R53.333

Total cost= R960+R53.33

           =R1013.33


ii)     Two tellers

1 teller R120 per hour

For two tellers for 8hours

R120*8*2 = R1920

 

Waiting

1 student =10min

40students = 40min

R80 waiting cost per hour

R80*40min/60min= R53.333/2

           = R26.67

Total cost = R1920+ R26.67

           = R 1946.67

 

iii)    I would choose one teller because the cost of operation is low.